OpenStudy (anonymous):

having promblem solving this linear equation: 2=-6/(x+4). Greatly appreciate some help?

OpenStudy (anonymous):

2= -6/(x+4) 2(x+4)=-6 2x+8=-6 2x=-14 x=-7

OpenStudy (anonymous):

to check 2=-6/(-7+4) 2=-6/-3 2=2 √

OpenStudy (anonymous):

what I don't understand, its in fraction form how did you move the den. (x+4). distribution law?

OpenStudy (anonymous):

Cross multiply

OpenStudy (anonymous):

Thank you, I have one more problem x-23=1/2(2x-46). since it's a fraction I don't know what to do.

OpenStudy (anonymous):

Need help solving this problem x-23=1/2(2x-46), the ans. is x-23=x-23. I worked inside the ( ) first by dividing by 2 which gave me x=-23. However, I don't know what to now with 1/2(x-23)?